What is B for A=1/2h(B+b)?
Answer by tutorcecilia(2105)

A=1/2h(B+b) [Solve for B]
A-b=1/2h(B+b-b)
A-b=1/2hB
(A-b)(2/h)=(2/h)(1/2h)B
(A-b)(2/h)=B
(2A-2b)/h=B
.
Check by plugging ((2A-2b)/h=B) back into the original equation.